23. (20 points) The following reaction was run at 285°C: COCl2(g) CO(g) + Cla(g) An equilibrium mixture of the three gases in i need help with question b
please solve b neatly
thank you for your time :)
Answer #1

The complete step wise calculation is given as below:-

6 It is given that, Ke=2.57X10 2 m =Equilibrium constant T= 28.5°C = (273.15 +28.5)k = 301.65K Reaction is given by:- cocla (Kp = kc (RT) An - 2 putting all the values in the above equation , kp= (2.57 X1o2m) (301.65KX0:0821 L.atm/malik) Kp= 2.57 * 1